OpenStudy (anonymous):

Solve the systems using substitution. 6x + 2y = 0 4x + y = -1

OpenStudy (luigi0210):

Solve for y in the second equation then plug that into the first equation.

OpenStudy (anonymous):

3x + 2y = -4 x + 2y = -8 5x - 4y = 14 x - 2y = -2 x + 3y = -4 4x + 5y = 5 x + 3y = -5 2x - y = 18 Would I do the same thing with these?? Solve for y and then plug it in?

OpenStudy (luigi0210):

Not all of them

OpenStudy (anonymous):

I don't want answers but could you tell me how I would do them?

OpenStudy (luigi0210):

Solve for a single variable.

OpenStudy (anonymous):

Gotcha and then just plug it in to find the other variable?

OpenStudy (luigi0210):

Yup, with practice it will be a breeze for you :)

OpenStudy (anonymous):

Thank you!

OpenStudy (luigi0210):

Anytime